OpenStudy (anonymous):

Geraldo has $15 less in his pocket than Zoe has. If Zoe had $7 more, she would have 4 times as much money as Geraldo. Let x = Geraldo's money and y = Zoe's money.

OpenStudy (anonymous):

x=G y=Z G=Z-15 Z+7=4G Z+7=4(Z-15) Z+7=4Z-60 3Z=67 Z=22.3 G=22.3-15=7.3 x=7.3 and y=22.3

OpenStudy (anonymous):

wrong 7.3*4 is not equal to 22.3

OpenStudy (anonymous):

but its "Z+7=4G" not "Z=4G" so I'm right